# GoLLIE : Annotation Guidelines improve Zero-Shot Information-Extraction

### 21 Feb 2024 v4

@@ Abstract
Large Language Models (LLMs) combined with instruction tuning have made significant progress when generalizing to unseen tasks. However, they have been less successful in Information Extraction (IE), lagging behind task-specific models. Typically, IE tasks are characterized by complex annotation guidelines which describe the task and give examples to humans. Previous attempts to leverage such information have failed, even with the largest models, as they are not able to follow the guidelines out-of-the-box. In this paper we propose GoLLIE (Guideline-following Large Language Model for IE), a model able to improve zero-shot results on unseen IE tasks by virtue of being fine-tuned to comply with annotation guidelines. Comprehensive evaluation empirically demonstrates that GoLLIE is able to generalize to and follow unseen guidelines, outperforming previous attempts at zero-shot information extraction. The ablation study shows that detailed guidelines is key for good results. Code, data and models are publicly available: https://github.com/hitz-zentroa/GoLLIE.
@@ 1 Introduction
The task of Information Extraction (IE) is highly challenging. This challenge is evident in the detailed guidelines, which feature granular definitions and numerous exceptions, that human annotators must follow to perform the task. The performance of current SoTA models heavily depends on the quantity of human-annotated data, as the model learns the guidelines from these examples. However, this performance significantly decreases when tested in new annotation schema (Liu et al., 2021a). The common practice in IE to achieve good results is to manually annotate in each new domain and schema from scratch, as almost no transfer exists across application domains. Unfortunately, this is unfeasible, both, in terms of financial cost and human effort.
Recent advancements in Large Language Models (LLM) (Min et al., 2023) have enabled the development of models capable of generalizing to unseen tasks. Thus, current zero-shot IE systems leverage the knowledge encoded in LLMs to annotate new examples (Sainz et al., 2022a; Wang et al., 2023a). As a by product of the pre-training process, models possess now a strong representation of what a person or an organization is. Therefore, they can be prompted to extract mentions to those categories from a text. However, this has a clear limitation: not every annotation schema111We define schema as the set of labels and their definitions. defines ”person” (or any other label) in the same way. For example, ACE05 (Walker et al., 2006) annotates pronouns as persons, while, CoNLL03 (Tjong Kim Sang & De Meulder, 2003) does not. IE tasks require more information than just label names, they require annotation guidelines.
Current LLMs have been trained to follow instructions, but they fail to follow annotation guidelines out-of-the-box. For instance, Figure 1 shows results on domain specific zero-shot Named Entity Recognition. The results of gpt-3.5-turbo when prompted with guidelines (Zhang et al., 2023a) are low, around 20 F1 score on Music or Politics domains. Building a system that enables high-performance zero-shot information extraction, reducing the dependence on costly human annotations, remains an open challenge.
@@ 2 Related Work
Large Language Models (LLMs) have made significant advancements toward the development of systems that can generalize to unseen tasks (Min et al., 2023). Radford et al. (2019) trained LLMs using a vast amount of internet data, finding that, pretrained models given natural language task descriptions can perform tasks such as question answering, machine translation or summarizing without explicit supervision. Building on this discovery, instruction tuning, often referred to as multitask fine-tuning, has emerged as the leading method to achieve generalization to unseen tasks. This process involves pre-training a model on a massive amount of unlabeled data and subsequently fine-tuning it on a diverse collection of tasks (Wang et al., 2022; Chung et al., 2022) phrased as text-to-text problems (Raffel et al., 2020). A natural language instruction or prompt is given to the model to identify the task it should solve (Schick & Schütze, 2021; Scao & Rush, 2021). Research has demonstrated that increasing the parameter count of the language model (Brown et al., 2020), coupled with improvements in the size and quality of the instruction tuning dataset, results in enhanced generalization capabilities (Chen et al., 2023; Zhang et al., 2022; Chowdhery et al., 2022; Muennighoff et al., 2023; Touvron et al., 2023a; b). LLMs have displayed impressive zero-shot generalization capabilities in various challenging tasks, including coding Wang & Komatsuzaki (2021); Black et al. (2022); Rozière et al. (2023), common sense reasoning Touvron et al. (2023a), and medical applications Singhal et al. (2023), among others.
@@
In the field of Information Extraction (IE), recent shared tasks (Fetahu et al., 2023) have shown that encoder-only language models such as XLM-RoBERTa (Conneau et al., 2020) and mDEBERTA (He et al., 2023) remain the most effective models. Attempts to utilize LLMs and natural language instructions for IE have been less successful (Tan et al., 2023; Zhou et al., 2023; Zhang et al., 2023a), as their performance lags behind that of encoder-only models. Before the billion parameters LLMs, indirectly supervised methods improve zero-shot IE by utilizing the knowledge learned from tasks like Textual Entailment (Sainz et al., 2021; 2022a; 2022b) and Question Answering (Levy et al., 2017). Obeidat et al. (2019) propose an entity typing method that encode label descriptions from Wikipedia as embeddings using an LSTM, which are then used to score the inputs. Methods that leveraged external knowledge were also successful on fine-grained zero-shot NER (Chen et al., 2021). Lu et al. (2022a) introduced a unified text-to-structure generation that can model different IE tasks universally. Lou et al. (2023) proposed converting IE tasks to a semantic matching problem, allowing their method to generalize to new domains and label ontologies not seen during training. Wang et al. (2023a) framed IE tasks as natural language descriptive instructions and trained an LLM across a diverse range of IE tasks. In evaluations on tasks with unseen label ontologies, their model outperformed other instruction-tuning methods.
@@ 3.1 Input-output representation
We have adopted a Python code-based representation (Wang et al., 2023b; Li et al., 2023) for both the input and output of the model. This approach not only offers a clear and human-readable structure but also addresses several challenges typically associated with natural language instructions. It enables the representation of any information extraction task under a unified format. The inputs can be automatically standardized using Python code formatters such as Black. The output is well-structured and parsing it is trivial. Furthermore, most current LLMs incorporate code in their pretraining datasets, indicating that these models are already familiar with this representation.
Figure 2 shows the three main parts of the format: schema definition, input text and output annotations. Schema definition forms the initial segment of the input. This section contains the information about the labels which are represented as Python classes; guidelines, articulated as docstrings; and representative annotation candidates presented in the form of code comments. The number of class definitions corresponds to the number of labels in the dataset. Classes are flexible and vary for each task. For example, classes for a NER dataset merely require an attribute to specify the text span that correspond to the class. On the other side, more complex tasks such as Event Argument Extraction (EAE) or Slot Filling (SF) demand more class attributes to categorize the task, such as a list of participants in an event (refer to examples in Appendix A). Input text is the second part of the input. The input text is represented as an string variable in Python. Output annotations is the part generated by the model. The model starts generating after result =. The annotations are represented as a list of instances of the classes defined on the schema definition part. Parsing the output is straightforward; executing the generated code in Python yields a list containing the result. This ease of parsing the output stands as a significant advantage of our model. A further detailed analysis of the efficiency of this approach is available in Appendix E.
@@ 3.2 Guidelines enhanced representation
The main contribution of this work is the use of the guidelines as part of the inference process to improve the zero-shot generalization. An example of a class definition with and without guidelines is shown in the Figure 3. Different datasets usually define guidelines on many different ways: some provides a complex definition of a label with several exceptions and special treatments and others just give a few representative candidates of the fillers of the label. To normalize the input format, we included the label definitions as class docstrings and the candidates as a comment for the principal argument (which is usually mention or span). Complex tasks such as EAE or SF requires additional definitions for the arguments or slots, to that end, we included small definitions as comments on each class argument.
@@ 3.3 Training regularization
We applied the following regularizations. Class order shuffling, for each example, the order of the input classes is randomly shuffled. This makes it more difficult for the model to memorize entire task definitions. Class dropout, we delete some of the input classes randomly. By eliminating few classes from both the input and output, we force the model to learn to only output instances of classes defined in the input. This not only encourages the model to focus on the schema definition but also minimizes the occurrence of hallucinations during inference. Guideline paraphrasing, we generate variations of the label definitions to prevent the model from easily memorizing them. We also think this will make the method more robust to different variations on the definition. Representative candidate sampling, similar to what we do with the paraphrases, for each input we sample 5 different candidates from a fixed pool of 10 per class. Class name masking involves substituting the label class names (e.g., Person) with placeholders, such as LABEL_1. This prevents the model from exploiting the label names during training, and forces it to attend and understand the guidelines.
@@ 4.1 Data
Evaluating zero-shot capabilities requires dividing the data into training and evaluation datasets. However, many benchmarks for Information Extraction are based on the same domain or share part of their schema. To ensure that the zero-shot evaluation is not affected by similar data, we have divided our set of benchmarks based on the domain of the data (a related topic is data contamination, that we discuss in Appendix G). For training we kept mostly datasets from News and Biomedical domains, for evaluation instead, we used datasets from diverse domains. This approach helps to avoid introducing any noise into the evaluation process. Among the evaluation datasets we included CrossNER (Liu et al., 2021b), a dataset that is split into many domains, for simplicity, we will call each domain as a separate dataset: AI, Literature, Music, Politics and Science. Also, we will call MIT Movie and MIT Restaurant as Movie and Restaurant. Table 1 contains the information about the data used in the experiments.
@@
We have trained the model to perform 5 different tasks: Named Entity Recognition (NER), Relation Extraction (RE), Event Extraction (EE), Event Argument Extraction (EAE) and Slot Filling (SF). However, we only evaluated the model on the three main tasks of interest: NER, EE and EAE. The other two tasks are added in the training data in order to add diversity and improve the flexibility of the model.
Few modifications has been done to two datasets in order to improve the quality of the model. First, the training data of Ontonotes 5 was reduced drastically as it was automatically annotated. Second, the TACRED dataset was converted from RE to SF in order to increase the complexity of the task. These modifications make our system not comparable with the state of the art on those tasks. However, our focus of interest is in the zero-shot evaluation and, therefore, the benefits (see Appendix A) are more interesting than adding 2 more comparable points on the supervised setup. In the CASIE dataset, we detected that the annotated event spans are inconsistent. The models typically annotate a sub-string rather than the entire span. Therefore, we evaluate all the models based on the predicted event categories, without considering the exact text span. For arguments, we use partial matching.
@@ 4.2 Language Models and Baselines
To train the models we use QLoRA (Hu et al., 2022; Dettmers et al., 2023). LoRA freezes the pre-trained model weights and injects trainable rank decomposition matrices into linear layers of the Transformer architecture. On a preliminary experiment this setup outperformed fine-tuning the entire model on the zero-shot tasks, while trained much faster (more details in Appendix D.4). We applied the LoRA to all linear transformer block layers as recommended by Dettmers et al. (2023). The models were trained for 3 epochs with an effective batch-size of 32 and a learning-rate of 3e-4 with cosine scheduler. Our training infrastructure was 2 NVIDIA’s A100 with 80gb each. More details about the training are given in the Appendix D.
@@
Our main point of comparison is Instruct-UIE (Wang et al., 2023a) as it is the approach closest to our system, but does not use guidelines. Another system considered for comparison is PromptNER (Zhang et al., 2023a), which proposes to prompt GPT-3.5 and T5 with definitions using Chain-of-Though in order to perform few-shot NER. Different from us, they did not fine-tune the model to attend to the guidelines. For fair comparison, we only considered the zero-shot results reported in the paper. In addition, other SoTA systems are added for comparison when results from Instruct-UIE and PromptNER are not available. Given that our systems is designed for the zero-shot scenario, the supervised experiments are intended to verify that our system does not degrade its performance. We thus we selected, for the supervised scenario, those systems among SoTA that share the most comparable setting with us.
@@ 5.1 Supervised evaluation
Compared to other systems our model achieves similar results in general. Focusing on the two datasets where our model under-performs significantly, WNUT and NCBIDisease, we find that task specific techniques are still needed. For instance, Wang et al. (2021) uses external knowledge to detect emergent and rare entities. In the NCBIDisisease dataset, models pre-trained on Biomedical domain corpora achieve best results (Kocaman & Talby, 2021). (Wang et al., 2023a) leverages Flan-T5, which has great proficiency on Biomedical domain tasks (Singhal et al., 2022). These improvements, however, are complementary to our proposal.
@@ 5.2 Zero-Shot evaluation
The results on the zero-shot are shown in Table 3. Overall, comparing to the baseline, the results are improved significantly when using guidelines on almost every dataset, with an absolute difference of 13 F1 points on average. Despite dividing the evaluation benchmarks based on the domain, there is always some overlap between labels of train and evaluation benchmarks. For instance, the datasets E3C and WikiEvents share a large part of their schema with datasets like BC5CDR, ACE05 and RAMS. This phenomena is reflected in the results.
GoLLIE surpass by a large margin the current zeri-shot SoTA methods Instruct-UIE (Wang et al., 2023a) and Entailment based IE (Sainz et al., 2022b). Compared to Instruct-UIE, the main differences are the backbone model, the amount of training data, and, the use or not of the guidelines. Instruct-UIE leverages the 11B FlanT5 which is a T5 fine-tuned on 473 NLP datasets. Respect to the data, Instruct-UIE leverages a total of 34 IE datasets (counting different tasks as datasets) from diverse domains, we only leverage 12 datasets. Contrary to our method they do not use guideline information. Still, our method performs significantly better suggesting that the guidelines have an important effect on the results.
PromptNER (Zhang et al., 2023a) also adds some definition information into the prompt in order to perform zero-shot NER. We compare our approach with them (represented as GPT-3.5) in Figure 1. Although their approach leverages guidelines too, our approach performs significantly better on all datasets, showing that LLMs (even with 175B parameters) struggle to follow guidelines. They solve this by adding examples in the context but still far behind on a comparable setting (T5-XXL).

### 06 Mar 2024 v5

@@ Abstract
Large Language Models (LLMs) combined with instruction tuning have made significant progress when generalizing to unseen tasks. However, they have been less successful in Information Extraction (IE), lagging behind task-specific models. Typically, IE tasks are characterized by complex annotation guidelines that describe the task and give examples to humans. Previous attempts to leverage such information have failed, even with the largest models, as they are not able to follow the guidelines out of the box. In this paper, we propose GoLLIE (Guideline-following Large Language Model for IE), a model able to improve zero-shot results on unseen IE tasks by virtue of being fine-tuned to comply with annotation guidelines. Comprehensive evaluation empirically demonstrates that GoLLIE is able to generalize to and follow unseen guidelines, outperforming previous attempts at zero-shot information extraction. The ablation study shows that detailed guidelines are key for good results. Code, data, and models are publicly available: https://github.com/hitz-zentroa/GoLLIE.
@@ 1 Introduction
The task of Information Extraction (IE) is highly challenging. This challenge is evident in the detailed guidelines, which feature granular definitions and numerous exceptions, that human annotators must follow to perform the task. The performance of current SoTA models heavily depends on the quantity of human-annotated data, as the model learns the guidelines from these examples. However, this performance significantly decreases when tested in new annotation schema (Liu et al., 2021a). The common practice in IE to achieve good results is to manually annotate each new domain and schema from scratch, as almost no transfer exists across application domains. Unfortunately, this is unfeasible, both, in terms of financial cost and human effort.
Recent advancements in Large Language Models (LLM) (Min et al., 2023) have enabled the development of models capable of generalizing to unseen tasks. Thus, current zero-shot IE systems leverage the knowledge encoded in LLMs to annotate new examples (Sainz et al., 2022a; Wang et al., 2023a). As a by-product of the pre-training process, models possess now a strong representation of what a person or an organization is. Therefore, they can be prompted to extract mentions of those categories from a text. However, this has a clear limitation: not every annotation schema111We define schema as the set of labels and their definitions. defines ”person” (or any other label) in the same way. For example, ACE05 (Walker et al., 2006) annotates pronouns as persons, while, CoNLL03 (Tjong Kim Sang & De Meulder, 2003) does not. IE tasks require more information than just label names, they require annotation guidelines.
Current LLMs have been trained to follow instructions, but they fail to follow annotation guidelines out of the box. For instance, Figure 1 shows results on domain-specific zero-shot Named Entity Recognition. The results of gpt-3.5-turbo when prompted with guidelines (Zhang et al., 2023a) are low, around 20 F1 score on Music or Politics domains. Building a system that enables high-performance zero-shot information extraction, reducing the dependence on costly human annotations, remains an open challenge.
@@
○ GoLLIE (Guideline-following Large Language Model for IE), an LLM fine-tuned to learn how to attend to the guidelines on a small set of well known IE tasks. Comprehensive zero-shot evaluation empirically demonstrates that GoLLIE outperforms the SoTA (Wang et al., 2023a) in zero-shot information extraction (see Figure 1).
@@ 2 Related Work
Large Language Models (LLMs) have made significant advancements toward the development of systems that can generalize to unseen tasks (Min et al., 2023). Radford et al. (2019) trained LLMs using a vast amount of internet data, finding that pre-trained models given natural language task descriptions can perform tasks such as question answering, machine translation, or summarizing without explicit supervision. Building on this discovery, instruction tuning, often referred to as multitask fine-tuning, has emerged as the leading method to achieve generalization to unseen tasks. This process involves pre-training a model on a massive amount of unlabeled data and subsequently fine-tuning it on a diverse collection of tasks (Wang et al., 2022; Chung et al., 2022) phrased as text-to-text problems (Raffel et al., 2020). A natural language instruction or prompt is given to the model to identify the task it should solve (Schick & Schütze, 2021; Scao & Rush, 2021). Research has demonstrated that increasing the parameter count of the language model (Brown et al., 2020), coupled with improvements in the size and quality of the instruction tuning dataset, results in enhanced generalization capabilities (Chen et al., 2023; Zhang et al., 2022; Chowdhery et al., 2022; Muennighoff et al., 2023; Touvron et al., 2023a; b). LLMs have displayed impressive zero-shot generalization capabilities in various challenging tasks, including coding Wang & Komatsuzaki (2021); Black et al. (2022); Rozière et al. (2023), common sense reasoning Touvron et al. (2023a), and medical applications Singhal et al. (2023), among others.
@@
In the field of Information Extraction (IE), recent shared tasks (Fetahu et al., 2023) have shown that encoder-only language models such as XLM-RoBERTa (Conneau et al., 2020) and mDEBERTA (He et al., 2023) remain the most effective models. Attempts to utilize LLMs and natural language instructions for IE have been less successful (Tan et al., 2023; Zhou et al., 2023; Zhang et al., 2023a), as their performance lags behind that of encoder-only models. Before the billion parameters LLMs, indirectly supervised methods improve zero-shot IE by utilizing the knowledge learned from tasks like Textual Entailment (Sainz et al., 2021; 2022a; 2022b) and Question Answering (Levy et al., 2017). Obeidat et al. (2019) propose an entity typing method that encodes label descriptions from Wikipedia as embeddings using an LSTM, which is then used to score the inputs. Methods that leveraged external knowledge were also successful on fine-grained zero-shot NER (Chen et al., 2021). Lu et al. (2022a) introduced a unified text-to-structure generation that can model different IE tasks universally. Lou et al. (2023) proposed converting IE tasks to a semantic matching problem, allowing their method to generalize to new domains and label ontologies not seen during training. Wang et al. (2023a) framed IE tasks as natural language descriptive instructions and trained an LLM across a diverse range of IE tasks. In evaluations on tasks with unseen label ontologies, their model outperformed other instruction-tuning methods.
@@ 3 Approach
○ GoLLIE forces the model to attend to the details in the guidelines, performing robustly on schemas not seen during training. On this section we deep dive into the details of our approach, describing how the input and output was represented and the regularization techniques used to force the model to attend to the guidelines.
@@ 3.1 Input-output representation
We have adopted a Python code-based representation (Wang et al., 2023b; Li et al., 2023) for both the input and output of the model. This approach not only offers a clear and human-readable structure but also addresses several challenges typically associated with natural language instructions. It enables the representation of any information extraction task under a unified format. The inputs can be automatically standardized using Python code formatters such as Black. The output is well-structured and parsing it is trivial. Furthermore, most current LLMs incorporate code in their pre-training datasets, indicating that these models are already familiar with this representation.
Figure 2 shows the three main parts of the format: schema definition, input text, and output annotations. Schema definition forms the initial segment of the input. This section contains information about the labels that are represented as Python classes; guidelines, articulated as docstrings; and representative annotation candidates presented in the form of code comments. The number of class definitions corresponds to the number of labels in the dataset. Classes are flexible and vary for each task. For example, classes for a NER dataset merely require an attribute to specify the text span that corresponds to the class. On the other side, more complex tasks such as Event Argument Extraction (EAE) or Slot Filling (SF) demand more class attributes to categorize the task, such as a list of participants in an event (refer to examples in Appendix A). Input text is the second part of the input. The input text is represented as a string variable in Python. Output annotations is the part generated by the model. The model starts generating after result =. The annotations are represented as a list of instances of the classes defined on the schema definition part. Parsing the output is straightforward; executing the generated code in Python yields a list containing the result. This ease of parsing the output stands as a significant advantage of our model. A further detailed analysis of the efficiency of this approach is available in Appendix E.
@@ 3.2 Guidelines enhanced representation
The main contribution of this work is the use of the guidelines as part of the inference process to improve the zero-shot generalization. An example of a class definition with and without guidelines is shown in Figure 3. Different datasets usually define guidelines in many different ways: some provide a complex definition of a label with several exceptions and special treatments and others just give a few representative candidates of the fillers of the label. To normalize the input format, we included the label definitions as class docstrings and the candidates as a comment for the principal argument (which is usually mention or span). Complex tasks such as EAE or SF require additional definitions for the arguments or slots, to that end, we included small definitions as comments on each class argument.
@@ 3.3 Training regularization
We applied the following regularizations. Class order shuffling, for each example, the order of the input classes is randomly shuffled. This makes it more difficult for the model to memorize entire task definitions. Class dropout, we delete some of the input classes randomly. By eliminating few classes from both the input and output, we force the model to learn to only output instances of classes defined in the input. This not only encourages the model to focus on the schema definition but also minimizes the occurrence of hallucinations during inference. Guideline paraphrasing, we generate variations of the label definitions to prevent the model from easily memorizing them. We also think this will make the method more robust to different variations on the definition. Representative candidate sampling, similar to what we do with the paraphrases, for each input we sample 5 different candidates from a fixed pool of 10 per class. Class name masking involves substituting the label class names (e.g., Person) with placeholders, such as LABEL_1. This prevents the model from exploiting the label names during training and forces it to attend and understand the guidelines.
@@ 4.1 Data
Evaluating zero-shot capabilities requires dividing the data into training and evaluation datasets. However, many benchmarks for Information Extraction are based on the same domain or share part of their schema. To ensure that the zero-shot evaluation is not affected by similar data, we have divided our set of benchmarks based on the domain of the data (a related topic is data contamination, which we discuss in Appendix G). For training we kept mostly datasets from News and Biomedical domains, for evaluation instead, we used datasets from diverse domains. This approach helps to avoid introducing any noise into the evaluation process. Among the evaluation datasets we included CrossNER (Liu et al., 2021b), a dataset that is split into many domains, for simplicity, we will call each domain as a separate dataset: AI, Literature, Music, Politics, and Science. Also, we will refer to MIT Movie and MIT Restaurant as Movie and Restaurant. Table 1 contains the information about the data used in the experiments.
@@
We have trained the model to perform 5 different tasks: Named Entity Recognition (NER), Relation Extraction (RE), Event Extraction (EE), Event Argument Extraction (EAE), and Slot Filling (SF). However, we only evaluated the model on the three main tasks of interest: NER, EE, and EAE. The other two tasks are added to the training data to add diversity and improve the flexibility of the model.
A few modifications have been made to two datasets to improve the quality of the model. First, the training data of Ontonotes 5 was reduced drastically as it was automatically annotated. Second, the TACRED dataset was converted from RE to SF to increase the complexity of the task. These modifications make our system not comparable with the state of the art on those tasks. However, our focus of interest is on the zero-shot evaluation and, therefore, the benefits (see Appendix A) are more interesting than adding 2 more comparable points on the supervised setup. In the CASIE dataset, we detected that the annotated event spans are inconsistent. The models typically annotate a sub-string rather than the entire span. Therefore, we evaluate all the models based on the predicted event categories, without considering the exact text span. For arguments, we use partial matching.
@@ 4.2 Language Models and Baselines
○ GoLLIE is a fine-tuned version of Code-LLaMA Rozière et al. (2023). Other backbone LLMs, such as LLaMA (Touvron et al., 2023a), LLaMA-2 Touvron et al. (2023b) or Falcon Penedo et al. (2023) were considered during the development, however, as our approach uses code to represent the input and output, Code-LLaMA model worked better on the preliminary experiments. In order to perform fair comparisons the baseline developed in this paper is based on Code-LLaMA as well. All the development of this paper was done with the 7B parameter version of Code-LLama, but, for a scaling analysis, we also trained the 13B and 34B parameter models.
@@
To train the models we use QLoRA (Hu et al., 2022; Dettmers et al., 2023). LoRA freezes the pre-trained model weights and injects trainable rank decomposition matrices into linear layers of the Transformer architecture. In a preliminary experiment, this setup outperformed fine-tuning the entire model on the zero-shot tasks, while training much faster (more details in Appendix D.4). We applied the LoRA to all linear transformer block layers as recommended by Dettmers et al. (2023). The models were trained for 3 epochs with an effective batch size of 32 and a learning rate of 3e-4 with a cosine scheduler. Our training infrastructure was 2 NVIDIA’s A100 with 80gb each. More details about the training are given in the Appendix D.
@@
Our main point of comparison is Instruct-UIE (Wang et al., 2023a) as it is the approach closest to our system, but does not use guidelines. Another system considered for comparison is PromptNER (Zhang et al., 2023a), which proposes to prompt GPT-3.5 and T5 with definitions using Chain-of-Though in order to perform few-shot NER. Different from us, they did not fine-tune the model to attend to the guidelines. For fair comparison, we only considered the zero-shot results reported in the paper. In addition, other SoTA systems are added for comparison when results from Instruct-UIE and PromptNER are not available. Given that our system is designed for the zero-shot scenario, the supervised experiments are intended to verify that our system does not degrade its performance. Thus we selected, for the supervised scenario, those systems among SoTA that share the most comparable setting with us.
@@ 5.1 Supervised evaluation
Compared to other systems our model achieves similar results in general. Focusing on the two datasets where our model under-performs significantly, WNUT and NCBIDisease, we find that task-specific techniques are still needed. For instance, Wang et al. (2021) uses external knowledge to detect emergent and rare entities. In the NCBIDisisease dataset, models pre-trained on Biomedical domain corpora achieve the best results (Kocaman & Talby, 2021). (Wang et al., 2023a) leverages Flan-T5, which has great proficiency on Biomedical domain tasks (Singhal et al., 2022). These improvements, however, are complementary to our proposal.
@@ 5.2 Zero-Shot evaluation
The results on the zero-shot are shown in Table 3. Overall, compared to the baseline, the results are improved significantly when using guidelines on almost every dataset, with an absolute difference of 13 F1 points on average. Despite dividing the evaluation benchmarks based on the domain, there is always some overlap between labels of train and evaluation benchmarks. For instance, the datasets E3C and WikiEvents share a large part of their schema with datasets like BC5CDR, ACE05, and RAMS. This phenomenon is reflected in the results.
GoLLIE surpasses by a large margin the current zero-shot SoTA methods Instruct-UIE (Wang et al., 2023a) and Entailment based IE (Sainz et al., 2022b). Compared to Instruct-UIE, the main differences are the backbone model, the amount of training data, and, the use or not of the guidelines. Instruct-UIE leverages the 11B FlanT5 which is a T5 fine-tuned on 473 NLP datasets. With respect to the data, Instruct-UIE leverages a total of 34 IE datasets (counting different tasks as datasets) from diverse domains, we only leverage 12 datasets. Contrary to our method they do not use guideline information. Still, our method performs significantly better suggesting that the guidelines have an important effect on the results.
PromptNER (Zhang et al., 2023a) also adds some definition information into the prompt in order to perform zero-shot NER. We compare our approach with them (represented as GPT-3.5) in Figure 1. Although their approach leverages guidelines too, our approach performs significantly better on all datasets, showing that LLMs (even with 175B parameters) struggle to follow guidelines. They solve this by adding examples in the context but are still far behind on a comparable setting (T5-XXL).